OpenStudy (anonymous):

solve algebraically... 10/x+3 + 10/3 = 6

OpenStudy (anonymous):

\[\frac{ 10 }{ x } +3\] or \[\frac{ 10 }{ x +3}\]

OpenStudy (anonymous):

\[10\div x + 3 + 10\div3 = 6\]

OpenStudy (anonymous):